OpenStudy (anonymous):

what is the factored forms of the following expressions 1. 6x^2+5x+1 2. 6x^2+17x+12 3. 12d^2+4d-1 4. 8g^2+6g-9 5.20x^2+22x-12

OpenStudy (lieutenantgeneral):

1.) x = -1/2 x = -1/3 2.) x = -3/2 x = -4/3 3.) 4.) g = -3/2 g = 3/4 5.) x = -3/2 x = 2/5

OpenStudy (anonymous):

thank you